### v1

# Multi-Parameter Molecular MRI Quantification using Physics-Informed Self-Supervised Learning

Abstract
Biophysical model fitting plays a key role in obtaining quantitative parameters from physiological signals and images. However, the model complexity for molecular magnetic resonance imaging (MRI) often translates into excessive computation time, which makes clinical use impractical. Here, we present a generic computational approach for solving the parameter extraction inverse problem posed by ordinary differential equation (ODE) modeling coupled with experimental measurement of the system dynamics. This is achieved by formulating a numerical ODE solver to function as a step-wise analytical one, thereby making it compatible with automatic differentiation-based optimization. This enables efficient gradient-based model fitting, and provides a new approach to parameter quantification based on self-supervised learning from a single data observation. The neural-network-based train-by-fit pipeline was used to quantify semisolid magnetization transfer (MT) and chemical exchange saturation transfer (CEST) amide proton exchange parameters in the human brain, in an in-vivo molecular MRI study (n=4). The entire pipeline of the first whole brain quantification was completed in 18.38.3 minutes, which is an order-of-magnitude faster than comparable alternatives. Reusing the single-subject-trained network for inference in new subjects took 1.00.2 s, to provide results in agreement with literature values and scan-specific fit results (Pearson’s r>0.98, p<0.0001).
Introduction
Magnetic resonance imaging (MRI) plays a central role in clinical diagnosis and neuroscience. This modality is highly versatile and can be selectively programmed to generate a large number of image contrasts[1], each sensitive to certain biophysical parameters of the tissue. In recent years, there has been extensive research into developing quantitative MRI (qMRI) methods that can provide reproducible measurements of magnetic tissue properties (such as: T1, T2, and T), while being agnostic to the scan site and the exact acquisition protocol used[2]. Classical qMRI quantifies each biophysical property separately[3], using repeated acquisition and gradual variation of a single acquisition parameter under steady state conditions. This is followed by fitting the model to an analytical solution of magnetization vector dynamics[4].
The exceedingly long acquisition times associated with the classical quantification pipeline have motivated the development of magnetic resonance fingerprinting (MRF)[5], which is an alternative paradigm for the joint extraction of multiple tissue parameter maps from a single pseudorandom pulse sequence. Since MRF data are acquired under non-steady state conditions[6], the corresponding magnetization vector can only be resolved numerically. This comes at the expense of the complexity of the inverse problem, namely finding tissue parameters that best reconstruct the signal according to the forward model of spin dynamics. Since model fitting under these conditions takes an impractically long time[7], MRF is commonly solved by dictionary matching, where a large number of simulated signal trajectories are compared to experimentally measured data[8]. Unfortunately, the size of the dictionary scales exponentially with the number of parameters (the "curse of dimensionality"[9]), which rapidly escalates the compute and memory demands of both generation and subsequent use of the dictionary for pattern matching-based inference.
Recently, various deep learning (DL)-based methods have been developed for replacing the lengthy dictionary matching with neural-network (NN)-based inference[10, 11, 12, 13]. While this approach greatly reduces the parameter quantification time, networks still need to be trained using a comprehensive dictionary of synthetic signals. Since dictionary generation may take days[12], it constitutes an obvious bottleneck for routine use of MRF, and reduces the possibilities for addressing a wide variety of clinical scenarios.
The complexity and time constraints associated with the MRF pipeline are drastically exacerbated for molecular imaging applications that involve a plurality of proton pools, such as chemical exchange saturation transfer (CEST) MRI[14]. While CEST has demonstrated great potential for dozens of biomedical applications[15, 16, 17, 18, 19, 20, 21], some on the verge of entering clinical practice[22], the inherently large number of tissue properties greatly complicate analysis[23]. This has prompted considerable efforts to transition from CEST-weighted imaging to fully quantitative mapping of proton exchange parameters [24, 25, 26, 27, 28]. Early CEST quantification used the fitting of the classical numerical model (based on the underlying Bloch-McConnell equations) after irradiation at various saturation pulse powers ()[27]. However, applying this approach in a pixelwise manner in-vivo is unrealistic because both the acquisition and reconstruction steps may require several hours. Later, faster approaches, such as quantification of the exchange by varying saturation power/time and Omega-plots [29, 30, 31, 32] still rely on steady-state (or close to steady state) conditions, and approximate analytical expressions of the signal as a function of the tissue parameters[33, 34]. Unfortunately, a closed-form analytical solution does not exist for most practical clinical CEST protocols, which utilize a train of off-resonant radiofrequency (RF) pulses saturating multiple interacting proton pools. Similarly to the quantification of water T1 and T2, incorporating the concepts of MRF into CEST studies provided new quantification capabilities[28, 35, 36, 37, 38] and subsequent biological insights, for example, in the detection of apoptosis after oncolytic virotherapy[12]. However, in order to make CEST MRF a realistic, robust, and attractive candidate for clinical use, the long dictionary generation time associated with each new application needs to be replaced by a rapid and flexible approach that adequately models multiple proton pools under saturation pulse trains.
Here, we describe a physics-based deep learning framework for rapid model fitting of the human brain tissue proton spin properties. While this approach is applicable for quantifying a variety of MRI parameters, we focus on a challenging CEST imaging scenario, involving multiple proton pools, a saturation pulse train, and non-steady-state MRF acquisition. The computational pipeline (Fig. 1) combines a spin physics simulator and a NN-based quantitative parameter reconstructor in a fully auto-differentiable manner[39]. Our system effectively solves and inverts the Bloch-McConnell ordinary differential equations (ODEs), which govern the multi-pool exchange, saturation, and relaxation dynamics of molecular MRI. Hence, we refer to this approach as "neural Bloch McConnell fitting" (NBMF). Importantly, the network can be be trained in a self-supervised manner, directly on the single-subject data of interest (inspired by related work on test-time-[40], internal-[41], and zero-shot-[42, 41, 43] learning). This circumvents the need for prior curation of a large training dataset, which is often inaccessible, especially for molecular MRI.
Results
In-vitro CEST quantification
A phantom composed of six vials with different combinations of L-arginine concentration and pH was assembled and scanned in a 3T clinical scanner (Prisma, Siemens Healthineers, Germany) using a previously published non-steady-state rapid CEST protocol[12, 44]. Excellent agreement was obtained between the NBMF-estimated and known L-arginine concentrations (Fig. 2a-b ): Pearson’s r = 0.986, p = , root mean square error (RMSE) = 8.4mM, mean absolute percentage error (MAPE) = 10.8%). The NBMF-reconstructed proton exchange rates were in excellent agreement with the corresponding values estimated by traditional MRF dictionary-matching (Fig. 2d-f): Pearson’s r = 0.999, p = , RMSE = 41.0 , MAPE = 13.2%. The pH dependence of the NBMF reconstructed exchange rate (Fig. 2d,e) was a good fit for a base-catalyzed proton exchange model (R2 = 0.94, p-value = ), as predicted by theory[45].
Quantifying the semisolid-MT and CEST proton exchange parameters in the human brain
The NBMF pipeline used in-vitro was modified for semisolid-MT and amide proton exchange parameter mapping in the human brain. Two 3D and rapid acquisition pulse sequences were applied, as described previously[12, 44]. The first sequence varied the saturation pulse frequency offset between 6-14 ppm (designed to encode semisolid-MT information), whereas the second sequence fixed it at 3.5 ppm (for amide proton parameter encoding). In both cases, a total of 31 raw information encoding images were generated, with the saturation pulse power randomly varied between 0-4 T[12, 45]. Water relaxation (T1 and T2) and field maps (B1 and B0) were acquired separately and used as an extra input to the neural reconstructor described in Fig. 1 in order to avoid water-pool- and inhomogeneity-associated biases, respectively[12]. A two-step NBMF (see Methods section) was used to fit the semisolid-MT and amide proton exchange parameters to the raw data.
Quantitative semisolid-MT and amide proton exchange parameter maps derived from a representative healthy volunteer are presented in Fig. 3 and Fig. 4, respectively. The resulting proton volume fractions and exchange rates were in agreement with the literature (although the large variability in previous reports is noted; see Fig. 3d-e and Fig. 4d-e). The mean values obtained for white/gray matter (WM/GM) were: for white matter and for gray matter.
The joint fit and training of the NBMF produced a pair of neural reconstructors, optimized on a single subject. We then re-applied the trained reconstructions to additional subjects in a fast inference mode. A representative example comparing the single-subject NBMF with rapid reconstructor reuse (following training on a different subject) is presented in Fig. 5. There is a good agreement between the results of the quantitative parameters obtained by both approaches, quantified by high correlation (Pearson’s r>0.98, p 0.0001), as well as satisfactory measures of structural similarity index and mean average percent errors (SSIM>0.8, MAPE<8%) for the vast majority of parameters/slices (Fig. 5e).
| BM numerical fitting |
Dictionary-based
pattern matching |
Supervised-learning NN-based mapping |
NBMF
(this work) |
| (A) Preparatory steps using synthetic data | none |
Hours to days
e.g., 62.5h (x50 nodes) for 70M-entries dictionary generation[12] |
Hours
e.g., ~4.5h for 60K-entries dictionary generation and network training[38] |
none |
| (B) First subject tissue parameters quantification |
Hours to days
e.g., 6h for reconstructing a pixel slice[48] |
Hours to days
e.g., 2.31h for reconstructing a pixel slice[12] |
~1 sec[12, 38]
(NN-based inference) |
<30 min
(joint fit & train) |
| (A+B) Total process till 1st subject maps completed | Hours to days | Hours to days | Hours | <30min |
| (C) subject mapping (N>1) | Hours to days | Hours to days |
~1 sec[12, 38]
(NN-based inference) |
~1 sec
(NN inference) |
Computational complexity and timing
The combined NBMF training and fitting procedure for all four semisolid-MT and amide proton volume fraction and exchange rate parameter maps from the whole brain of a single subject took minutes on a standard GPU-equipped (GeForce RTX 3060) desktop workstation, of which, the two-pool quantification of the semisolid MT pool parameters took min. Re-applying the trained reconstructors for whole-brain parameter mapping on a new subject took sec.
Table 1 compares the computational time required for NBMF and alternative parameter fitting/quantification approaches. Overall, the complete quantification pipeline (preparation included) takes less than 30 min for NBMF, compared to at least several hours for traditional Bloch-Fitting or MRF-based dictionary generation and use for matching or supervised NN training.
Discussion
Automatic differentiation of the Bloch-McConnell (BM) equation solutions.
Quantification of semisolid MT/CEST proton exchange parameters under non-steady-state conditions is a notable example of a biophysical estimation where the forward model is perceived as too complex for direct inverse solution via fitting (requiring several days for a single whole brain reconstruction[48]). While the solution can be approximated via MRF, the large simulated signal dictionaries[10, 12] associated with multi-pool imaging also demand significant computational resources[50, 51], limiting their routine use and the development of new pulse sequences. A recently reported dictionary-free alternative[52] proposed unsupervised learning for semisolid-MT parameter quantification. However, this method assumes continuous pulse irradiation, which is not available in many clinical scanners, and also relies on analytical solutions, which are not compatible with multi-pool pulsed CEST imaging.
The dictionary-free method presented in this work overcomes all such limitations. Our approach is based on a fundamental insight: by proper formulation, ODE models considered only numerically solvable can become step-wise analytical, and thereby compatible with automatic differentiation-based optimization. Specifically, the suggested formulation enables GPU-bound matrix inversion and exponentiation, which translates into efficient gradient descent via back-propagation. Combining this concept with a recently reported high-performing automatic differentiator[39] provides a new option for solving complex biophysical estimation tasks such as pulsed CEST quantification, demonstrated here. Compared to standard model fitting, this approach avoids (i) computationally expensive and inaccurate purely-numerical derivatives computed via multiple evaluations, and (ii) explicit analytical approximations, which can only be applied to a limited sub-cohort of cases and lack generalization (e.g., unsuitable for a multi-proton-pool pulsed-RF saturation). Unlike dictionary-based[45] or reconstruction network-based MRF[28], the suggested approach can provide parameter estimates that allow the model to best describe the raw data.
The gradient of the forward model can be directly used for simple fitting of the unknown ODE coefficients corresponding to the physical parameters of interest (see voxelwise BMF in the Methods section and Supplementary Section 3). However, when the core forward model automatic differentiator is also integrated into a self-supervised learning pipeline (NBMF, Fig. 1), a joint neural representation of the signal-to-parameter transformation can be trained and stored with little extra computational cost. This enables: (i) Faster convergence, which scales well with the number of voxels up to the full brain size, leveraging redundancy towards a spatially smoother solution (Supplementary Figure 2). (ii) Later reuse for real-time inference on new subjects within a similar imaging scenario.
The human brain imaging results (Figs. 3-5) reveal the potential for using an autodiff-compatible Bloch-McConnell solver for parameter quantification, while training a simple multilayer perceptron (MLP) voxelwise. Combined with a 3D whole brain acquisition routine (which rapidly generates hundreds of thousands of voxels), the suggested system provides a method for rapid and efficient single-subject learning. Future work could study different NN architectures with spatial awareness (via convolutional or attention layers), as well as larger datasets composed of multiple subjects and/or synthetic data.
The proposed approach could be further exploited for other tasks across the CEST-MRF pipeline, such as accelerated dictionary synthesis[53, 54] and pulsed-wave irradiation compatible CEST protocol discovery and optimization[51, 50, 55, 56, 57]. Furthermore, NBMF is directly applicable to anatomical-MRF (proton density, T1, T2) dictionary-free parameter quantification and conversely, to classical non-MRF molecular MRI, such as pulsed multi-B1 Z-spectra fitting[58, 29] (see Supplementary Section 1). While auto-differentiation of the Bloch equations has previously been leveraged for several MRI-related applications[59, 51, 56, 60, 61], to the best of our knowledge this is the first report of utilizing this concept for a generalized Bloch-McConnell-fitting task. Beyond molecular MRI and MRF, this approach can also be applied to any other diagnostic and biophysical domains that involve dictionary-matching[62].
Learning to estimate ordinary differential equation (ODE) coefficients from observed data.
The general strategy underlying NBMF can be applied to any inverse problem that involves fitting ODEs to observations of a dynamical system. In the biomedical realm alone, this includes cardiovascular[63, 64], pharmacokinetic[65, 66], and epidemiological[67, 68] modeling, among many other tasks. In parallel to the exponential growth and improvement in AI performance, the last decade has witnessed a surge of interest in harnessing deep learning for physics-based problem solving. These efforts have most often been directed into two routes: (i) seeking a solution to a partial differential equation (PDE) as an output of a physics-informed neural network (PINN) that operates on spatial and temporal coordinates[69, 70, 71, 72, 73]; widely applied for spatially-resolved dynamics in solid[74]- and fluid[75] mechanics, heat transfer[76], power systems[77], weather/climate[78], and diffusion MRI[79]. (ii) Modeling parts of the equation with a NN, yielding a neural ODE/PDE[80], often employed as a semi-parametric approach for model discovery[81, 82, 83]. Interestingly, the relatively simple direct inverse solution approach described here (Fig. 1), whereby a NN is trained to infer the coefficients of an ODE model from a few samples of the dynamics, has not received similar attention. This could open opportunities for the current work to inform new approaches to ODE-driven inverse problems across a multitude of tasks.
Conclusions
The NBMF framework enables rapid, dictionary-free, pulsed-saturation- and multiple proton-pool-compatible semisolid MT/CEST-MRF quantification. By combining a GPU-accelerated auto-differentiable numerical ODE solver and self-supervised deep learning, the NBMF pipeline is able to match alternative AI-reconstruction based inference, while removing the need for dictionary synthesis and reconstruction network pre-training. NBMF is three orders of magnitude faster than traditional Bloch fitting, and provides a one-stop-shop for reconstructing quantitative molecular MRI data. The underlying approach has potential applications across a wide variety of ODE-driven inverse problem tasks.
Methods
CEST Phantoms
A set of six 10 ml L-arginine (L-arg, chemical shift = 3 ppm, Sigma-Aldrich) phantoms was prepared at a concentration of 25, 50, or 100 mM. The phantoms were titrated to different pH levels between 4.0-5.0 and placed in a 120 mm diameter cylindrical holder (MultiSample 120E, Gold Standard Phantoms, UK), filled with saline.
Human subjects
Four healthy volunteers (three females/one male, with average age 23.750.83) were scanned at Tel Aviv University (TAU), using a 3T MRI equipped with a 64-channel coil (Prisma, Siemens Healthineers). The research protocol was approved by the TAU Institutional Ethics Board (study no. 2640007572-2) and the Chaim Sheba Medical Center Ethics Committee (0621-23-SMC). All subjects gave written, informed consent before the study.
MRI Acquisition
All acquisition schedules were implemented using the Pulseq prototyping framework[84] and the open-source Pulseq-CEST sequence standard[85]. The MRF acquisition protocols were implemented as described previously[12, 44], with an unsaturated M0 image added at the beginning of each sequence. A spin lock saturation train (13 100 ms, 50% duty-cycle) was used for each one of the 30 additional iterations of the sequence, which varied the saturation pulse power between 0 and 4 T (average pulse amplitude). The saturation pulse frequency offset was fixed at 3 ppm for L-arginine phantom imaging[45], 3.5 ppm for amide brain imaging[12], or varied between 6 and 14 ppm for semisolid MT imaging[44]. The saturation block was followed by a 3D centric reordered EPI readout module[86, 87], providing a 1.8 mm isotropic resolution. The full sequences can be accurately reproduced using previously published Pulseq (.seq) files[53]. Each 3D MRF acquisition took 2:36 (min:s). The same readout module was used for acquiring additional B0, B1, T1, and T2 maps, via WASABI[88], saturation recovery, and multi-echo sequences, respectively. The total scan time per subject was 9 min.
MRI data pre-processing
In vitro images with no L-arginine vials, partial vials, or severe image artifacts were removed. Regions of interest (ROIs) were defined using circular masks. In-vivo brain images were motion-corrected and registered using elastix[89]. WM/GM ROI segmentation of the T1 map was performed using statistical parameter mapping (SPM)[90]. Quantitative reference CEST-MRF values (Fig. 2) were obtained using dot-product matching, as extensively described previously[45, 53].
NBMF architecture for semisolid-MT and CEST quantification
The self-supervised learning framework comprises two main components (Fig. 1, Top):
(A) Reconstructor - a fully-connected multi-layer perceptron (MLP) NN, applied voxel-wise on the raw input data[10, 12, 13, 91]. The NN is composed of three layers, with 256 neurons and ReLU activation in each hidden layer. The output layer consists of 5 neurons, encoding the estimates for the proton volume fraction and exchange rate of the compound of interest and their joint uncertainty expressed as noise covariance. It utilizes a sigmoid activation, with the output scaled to a predefined range of the parameter values, which effectively defines the prediction boundaries as follows: semisolid proton volume fraction f[0, 30] (%), semisolid proton exchange rate k (s-1), amide proton volume fraction f[0, 1.2] (%), amide proton exchange rate k[0, 400] (s-1), L-arginine concentration [L-arg] (mM) and Nα-amine (of L-arginine) proton exchange rate k[100, 1400] (s-1)[45]. Several auxiliary maps X, including water relaxation T1, T2, and B0/B1 inhomogeneities, are appended to the MRF raw data D to be used as inputs for the tissue parameter estimation: where are the weights to be trained.
(B) Simulator - a differentiable multi-pool spin physics solver. A numerical simulation of the piece-wise-constant coefficient Bloch-McConnell (BM) differential equations was implemented in the open-source JAX[39] computing framework, leveraging its strong auto-differentiation and GPU-acceleration capabilities for matrix operations. The simulator concatenated and chained the calculations of the BM closed-form solution across all pulses and delays of the protocol. This was carried out by inversion and exponentiation of the 9 9 BM-matrix , which expresses all precession, saturation, relaxation and exchange terms of the multi-pool magnetization vector () dynamics, as previously defined[34]:
This solver is compatible with the rectangular pulse-train shape employed in this study and others[12, 44, 85], while arbitrary pulse shapes can be supported using a simple matched-RMS approximation, or to any order through a Magnus expansion[92]. For two-pool imaging cases (such as semisolid MT data acquired using saturation pulses with a frequency offset higher than 6 ppm), additional acceleration was obtained by implementing the interleaved saturation–relaxation (ISAR2) approximate analytical solution[34] of the saturation stage. The RF pulses of spin-lock and readout flips were approximated as hard pulses generating precise flip angle rotations.
The model is designed to represent the whole sequence by simulating the Z-magnetization dynamics during the recovery, saturation, and readout stages, provided that spoilers are applied. For each of the two (semisolid MT/amide) sequences, the Nx31 non-steady-state MRF measurements from approximately 180K brain voxels were normalized using an unsaturated M0 reference image. Thus, the resulting acquired data is directly related to the magnetization vector governed by Eq. (1), at the end of the saturation pulse block. Therefore, given , an estimate of the sought parameters, the simulator provides a re-synthesis of the data as: ) , where and B1 are the saturation pulse frequency offsets and powers implemented in the MRF protocol, and X are any known tissue parameters.
The NBMF reconstruction of the semisolid-MT proton exchange parameters from the first (1) subject, was obtained by using the MT-MRF data alongside independently quantified auxiliary parameter maps , for training the weights of a neural reconstructor , designed to quantify the associated proton exchange parameters (). To that end, the NBMF optimizes the following self-supervised objective of consistency with the biophysical model :
The norm was used as the regression loss. A cosine-decay learning rate schedule and simple early-stopping upon convergence (loss trend reaching plateau) were applied. Augmentation by noise was applied, twice: (a) Adding a 0.1% uniform noise to the raw samples (b) Adding a Gaussian noise to the tissue parameters estimate, using covariance derived from extra outputs of the NN, inspired by a recent work[93].
This process was repeated using the non-steady-state amide raw MRF data for NBMF quantification of the amide proton exchange parameters (). For human brain experiments, we also appended the semisolid MT pool parameter estimates and (obtained from the semisolid MT NBMF procedure) to the auxiliary vector . This vector served as input for the amide reconstructor and the 3-pool biophysical model , so that: [12]. For the two-pool L-arginine phantom experiments, the auxiliary parameters were assigned constant values based on previous reports (T1w = 2800 ms, T2w=1200 ms)[44].
Importantly, we obtain both the subject-specific proton exchange parameters and the trained reconstructor at the convergence of the NBMF. This enables ultra-fast quantification of the proton exchange parameters from a new subject (2) (Fig. 1 bottom). Notably, this rapid inference is only applicable for new data drawn from the same distribution and cannot be applied to entirely new systems (such as muscle creatine quantification using brain-data trained NBMF), although NBMF could be retrained for a new imaging scenario of interest.
As a natural ablation of the system by removing the neural component, the auto-diff simulator can be used for direct voxelwise parameter fitting: , referred to here as voxelwise Bloch-McConnell fitting (VBMF). This simpler process can be described in the context of Fig. 1 as stopping the gradients at the tissue parameters, which now assume the role of independent per-voxel variables. Apart from the obvious drawback of not yielding a neural reconstructor, VBMF’s performance is inferior to NBMF for brain imaging (Supplementary Fig. 2), establishing the need for the full NBMF. However, it is a viable direct method for in vitro analysis that gives equivalent or even slightly better results, for that small data.
Finally, additional acceleration was achieved by parallelization of the computational graph across consecutive readout pairs , decoupling the single-iteration simulators . Assuming that the snapshot captures the preceding spin history evolution, the re-synthesis stage is now formulated as and embedded in Eq. (2) as such. See Supplementary Section 1 for further elaboration.
Statistical analysis
The structural similarity index measure was calculated using the open-source SciPy scientific computing library for Python. In slice-statistic box plots (Fig. 5e), the central horizontal lines represent median values, box size represents the two central (2nd, 3rd) quartiles, whiskers represent 1.5 the interquartile range above and below the upper and lower quartiles, and circles represent outliers. In the voxel-statistic box plots (Figs. 3, 4) the whiskers represent the 90 central percentiles. Numerical results in the text are presented as meanSD.
Data and Code availability
A dedicated repository containing the associated code and sample data will be made public upon acceptance.
Acknowledgements
This project was funded by the European Union (ERC, BabyMagnet, project no. 101115639). Neither the European Union nor the granting authority can be held responsible for views and opinions expressed, which are those of the authors alone. This project received funding from the Ministry of Innovation, Science and Technology, Israel, and a grant from the Tel Aviv University Center for AI and Data Science (TAD, The Blavatnik AI and Data Science Fund).
Author contributions statement
A.F. and O.P. conceived the computational framework with N.V. providing valuable input. M.Z., N.V. and O.P. developed and/or conducted the imaging protocols and data acquisition. A.F. was responsible for the AI design, optimization, and data analysis. A.F. wrote the manuscript. O.P. supervised the project. All authors reviewed the manuscript.
Competing Interests
The authors declare no competing interests.
Supplementary section 1: Extending NBMF for Z-spectra parameter fitting
The main focus of the manuscript is on fitting non-steady-state pulsed-wave (PW) CEST-MRF data. However, both CEST-MRF and Z-spectra can be unified and represented under a generalized CEST protocol[29]. In both cases the same blocks (readout, recovery, and saturation) are applied, with the saturation pulse power and/or frequency offset varying between different images. The recently adopted accelerated Z-spectra collection[29] is also non-steady-state, with larger but still finite recovery times (e.g., = 3 sec); the effect is approximated[29] by a history-independent constant factor (i.e., assume saturation and acquisition bring z-magnetization to zero, so the pre-saturation magnetization obeys: , e.g., 0.95 for 1 sec). In Z-spectra modeling this is typically joined by other approximations for the complicating effects of multiple pools, pulsed and finite (partial saturation) RF irradiation. Here, we analyze Z-spectra sequences as an instance of CEST-MRF, with the saturation parameters following a specific sequential pattern, without the above approximations for the dynamics.
The insight we offer here is that the initial condition of each experiment in each voxel is known (to the precision permitted by the measurement) by virtue of the previous readout. This removes the need for a sequential simulation of the entire magnetization history. The initial (normalized) pre-saturation magnetization can be estimated for each iteration by processing the result of the "snapshot"[94] acquisition through simulation of all flips and relaxations throughout the readout and recovery stages as follows:
where represents the operator of (normalized) Z-magnetization evolution across the readout stage involving multiple small-flip-angle rotation and pause blocks, with spoilers applied. It is assumed the acquired decoded signal can be translated to Z-value via normalizing voxel-wise by the equilibrium (no-saturation) signal reference, e.g. the first B1[n=0]=0 iteration. The prediction for the next acquisition will thus be:
where is a Bloch-McConnell simulator of the saturation stage, is the overall evolution across one iteration, including readout (), recovery (as formulated explicitly in Eq. (1)) and saturation () stages, and .
This enables us to avoid sequential modeling of the hundreds of iterations that constitute the multi-B1 Z-spectra collection, and to achieve parallel processing for both forward simulation and inverse fitting. We thus treat both types of experiments the same, by ignoring the Z-spectra special sequence and viewing it just like CEST MRF: a batch of magnetization evolution instances with known initial and final conditions, varying saturation RF parameters and common unknown tissue parameters. These values can be obtained by fitting, i.e. seeking best reconstruction of final (normalized) magnetizations from initial ones under the model (Eq.(4)), computed as the regression: .
Within the technical context of the NBMF framework, the above analysis leads to two options: sequential or parallel execution of the simulation. The parallel option results in faster training in the full numerical (3-pool) case, especially for Z-spectra, which involve an order-of-magnitude longer sequence. We therefore use this option for applications of NBMF to quantification of parameters from either CEST-MRF (see main text) or multi-B1 Z-spectra (see next section).
More generally, the insight regarding the ability to use non-constant computed from the previous iteration can be applied to various existing analytical or numerical methods for fitting Z-spectra. Furthermore, it can inform acquisition design by suggesting reduced or possibly non-uniform recovery times.
Supplementary section 2: Z-spectrum fitting speed benchmark
The speed of the proposed approach was compared to traditional numerical BM-fitting[29, 95], as implemented using a common CEST MATLAB repository[29] (https://github.com/cest-sources/BM_sim_fit, commit 029e00c). The same desktop workstation, equipped with 24 Intel i9 CPUs and a single Nvidia GeForce RTX 3060 GPU was used in both cases. The benchmark data were obtained by simulating synthetic multi-B1 Z-spectra[29], assuming three proton pools: (a) water, (b) amide, and (c) semi-solid (MT) with parameters oriented for white matter (T1=1.5s, T2w=49ms, T2s=40ms, T2ss=40s, fs=0.2%, ksw=300, fss=20%, kssw=70). The field strength was set to B0=7T, the saturation pulse powers were B1=[.6, .9, 1.5, 2.0] T, and the saturation pulse frequency offset was varied non-uniformly between -30 and 30 ppm, with the majority of samples concentrated between -5 and 5 ppm in 0.25 ppm increments, for a total of 61 samples per B1 value and 244 overall. A 1% noise was added to the signal. The full sequence details are available in: https://github.com/cest-sources/MultiB0_B1_qCEST_brain. An excellent fit (=99.8) was obtained by both methods (see Supplementary Fig. 1). The process was repeated using different starting values for the proton volume fraction and exchange rate estimates, demonstrating similar convergence to the correct values in all cases (Supplementary Fig. 1c,d), even for the more challenging ksw estimate.
The speed performance of both methods is presented in the Supplementary Table 1. Traditional BM fitting requires 40 seconds per voxel, with the overall time scaling linearly with the number of voxels to reach a total reconstruction time of 160 hours, impractical for 3D human imaging. Initializing NBMF processing in the JAX framework[39] requires roughly a minute for graph compilation and GPU allocation, but this is quickly amortized by the speed of subsequent highly optimized rapid fitting iterations. For a medium-sized single brain image slice of 128x12816K voxels, the runtime for NBMF is 3 orders of magnitude shorter than the benchmark.
| Numerical fitting method | NBMF (this work) | BM fitting[29] |
| 10 voxels reconstruction time | 30 sec (measured) | 400 sec (measured) |
| 16K voxels reconstruction time | 200 sec (measured) | 160 hours (estimated) |
Supplementary section 3: VBMF discussion and comparison with NBMF
Here we expand on the gradient-based voxelwise Bloch-McConnell fitting (VBMF), a natural simplification of the NBMF pipeline, achieved by stripping the neural component and stopping the gradient at the voxelwise tissue parameters which thereby become the trainable "variables" (see Fig. 1 and Methods section of the main text). Thus, the VBMF is directly comparable to legacy Bloch fitting of multi-B1 Z-spectra (or CEST-MRF acquired data, given the parallel mode, see Supplementary Section 1). Note that as a first-order optimizer, its convergence is theoretically inferior to second-order options such as Levenberg-Marquardt (LM) which is considered standard (e.g., BM_sim_fit[29] employs LM implicitly through matlab’s lsqcurvefit defaults). In practice, however, the machine-analytic GPU-accelerated solution ensures very fast convergence, as demonstrated in the preceding section. Supplementary Fig. 2 presents a comparison of tissue parameter maps reconstructed with NBMF and VBMF from a semisolid MT MRF data acquired in a healthy human volunteer. Similar results are obtained for the estimates of the concentrations, with a denoising effect for the exchange rate observed for NBMF. This a known attribute of neural networks. While neighboring voxels can have significantly different fit results by traditional BM fitting, NBMF, on the other hand, fits a single (shared between pixels) model, which improves data utilization and robustness to signal deviations. Notably, these benefits occur despite the network architecture being a trivial voxelwise MLP, which does not learn the brain anatomy. We predict that incorporation of spatial learning via the use of convolutional neural networks (CNNs), potentially combined with training on multiple subjects, will constitute a promising future direction for research.
Thus, VBMF might be a viable accelerated alternative to existing fitting methods, broadly applicable to generalized Z-spectra/CEST-MRF in-vitro experiments. Like other direct fitting methods, VBMF offers conceptual simplicity and guarantees the best fit or minimized discrepancy per voxel. This characteristic, shared with dictionary-matching methods (up to the precision limits imposed by grid quantization), is not retained in faster neural-inference methods.
However, NBMF is preferable for human brain imaging. As emphasized in the main text, it is able to provide a neural network (NN) reconstructor enabling real-time inference on new subjects at no extra cost. This is because the highly optimized GPU implementation makes the added computational burden of the small NN negligible. In addition, NBMF reaches convergence faster, possibly within just a few epochs, because the data redundancy relieves the model from the need to see each individual voxel many times.
Supplementary section 4: NBMF results for the estimation of quantitative exchange parameters from semisolid MT-MRF and CEST-MRF data of n = 4 healthy volunteers
Supplementary section 5: Auxiliary data overview